Circular No. 8310

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004aw IN NGC 3997 Independent discovery reports have been received of an apparent super-

nova in NGC 3997 from T. Boles (Coddenham, England, 0.35-m reflector) and from K. Itagaki (Teppo-cho, Yamagata, Japan, 0.60-m reflector; via S. Nakano, Sumoto). Boles measures the position of SN 2004aw as α = 11^{h}57^{m}50^{s}.24, δ = +25^{o}15′55′′.1, which is ≈ 27′′.7 east and 19′′.8 south of the center of NGC 3997. Itagaki provides position end figures 50^{s}.26, 55′′.1. Unfiltered CCD magnitude estimates of SN 2004aw: 2001 Nov. 25 UT, [18.5 (Itagaki); 2003 May 7, [19.5 (Boles); 2004 Mar. 13, [18.5 (Itagaki); 19.851, 17.1 (Boles); 20.51, 15.8 (Itagaki); 20.646, 16.2 (R. and Y. Kushida, Yatsugatake South Base Observatory). Boles adds that the new object is not present on 1996 Digitized Sky Survey red and blue images.

V5114 SAGITTARII Further to IAUC 8307, H. Yamaoka, Kyushu University, notes that

V5114 Sgr was also not visible on an ASAS-3 image taken on Mar. 12.38 UT (limiting mag ∼ 13.5). He adds that a CCD image taken by K. Itagaki on Mar. 18.8 shows the nova at α = 18^{h}19^{m}32^{s}.19, δ = −28^{o}36′35′′.0 (equinox 2000.0; ∼ 70 UCAC-2 reference stars, uncertainty ∼ 0′′.3). Nothing is visible on a Digitized Sky Survey red image within 2′′ of this position.

B. E. Schaefer, Louisiana State University, writes that CCD images obtained with the 0.8-m telescope at McDonald Observatory show no visible light echo. Reported CCD magnitudes and colors for V5114 Sgr: Mar. 19.336 UT, V = 8.64 ± 0.03, B − V = +0.66 ± 0.05 (W. Liller, Viña del Mar, Chile); 19.467, V = 8.76 (J. D. West, Mulvane, KS); 19.471, B = 9.28 (West); 19.49, U = 8.87, B = 9.23, V = 8.65, R = 8.01, I = 7.38 (Schaefer); 20.49, U = 9.08, B = 9.53, V = 9.00, R = 8.19, I = 7.42 (Schaefer); 19.695, V = 8.73, U − B = −0.50, B − V = +0.51, V − R = +0.66, V − I = +1.18 (A. C. Gilmore, Mt. John Observatory; uncertainties ∼ 0.02 mag).

2004 FH G. W. Christie, Auckland Observatory, reports that his unfiltered CCD

photometry (3-s exposures approximating the R band) of 2004 FH (cf. MPEC 2004-F24, 2004-F26) during Mar. 18.487–18.499 shows light variations with a period of ≈ 90 s (nine clear maxima) and a variable amplitude of ≈ 0.9 magnitude (ranging from mag 13.8 to 12.9).

2004 March 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green